A number of identical cells, n, each of emf E, internal resistance r connected in series are Charged by a d.c. source of emf E¢, using a resistor R.
(i) Draw the circuit arrangement.
(ii) Deduce the expressions for (a) the charging current and (b) the potential difference across the combination of the cells.
 

Grade:12

1 Answers

Arun
6 years ago
Resistance in the circuit R , emf of each cell is e and internal resistance is r, connected in series and total number of cell is n and the DC source emf  is E.
So the resistance of the circuit =nr+R and emf =ne+E
(i)Therefore current through the circuit to charge the cells is , I  = (E - ne)/(nr + R)
(ii) Potential difference across the combination of cells = ne